OpenStudy (anonymous):

Which of the following best describes the equation below? 2(x + 10) + 8x = 10x - 20 A. infinite real solutions B. no real solutions C. exactly one real solution, x = D. exactly one real solution, x = -1

OpenStudy (jeffyblood):

I think that it is B or A

OpenStudy (aripotta):

2x + 20 + 8x = 10x - 20 10x = 10x 0 = 0 when we substitute 2 for x: 2(2) + 20 + 8(2) = 10(2) - 20 4 + 20 + 16 = 20 - 20 40 = 0, which is obviously not true so no real solutions

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

If r = 6 units and h = 9 units, what is the volume of the cylinder shown above? Use 3.14 for . A. 1,526.04 cubic units B. 452.16 cubic units C. 339.12 cubic units D. 1,017.36 cubic units

OpenStudy (jeffyblood):

do you have a picture of the cylinder?

OpenStudy (aripotta):

we don't really need the picture tho. we just need the info

OpenStudy (aripotta):

\[v~=~\pi r ^{2}h\]

OpenStudy (aripotta):

so v = 3.14(6^2)(9) do that and you'll have your answer

OpenStudy (anonymous):

can you draw the equation because im pretty lost. thats how dumb i am.

OpenStudy (jeffyblood):

Your not dumb just confused

OpenStudy (aripotta):

the formula for the volume of a right cylinder is pi times radius squared times height